Landmark, Inc. hired you as a consultant to help them estimate its cost of capital. You have been provided with the following data: its most recently paid dividend is $1.25; its current market price is $31.50; its ROE = 12% and its dividend payout ratio is 40%. New common stock will have an 8% flotation cost. Based on the DCF approach and the Retention Growth Model, what is the cost of equity from new common stock? Enter your answer rounded to two decimal places. Do not enter % in the answer box. For example, if your answer is 0.12345 or 12.345% then enter as 12.35 in the answer box.
Answer #1

growth Rate = ROE * (1- Div Pay Outratio )

= 12% ( 1 - 0.4)

= 12% * 0.6

= 7.2%

D1 = D0(1+g)

= 1.25 ( 1 +0.072)

= 1.25 * 1.072

= 1.3438

Ke = [ D1 / P0(1 - Floation cost ) ] + Growth Rate

= [ 1.3438 / 31.50 ( 1 - 0.08 ) ] + 0.072

= [ 1.3438 / 31.50 (0.92 ) ] + 0.072

= [ 1.3438 / 28.98 ] + 0.072

= 0.0464 + 0.072

= 0.1184 i.e 11.84%